A disk between vertebrae in the spine is subjected to a shearing force of 600N. Find its shear deformation using the shear modulus of 1.0 times 10raised to 9 Pa. The disk is equivalent to a solid cylinder 0.7 cm high and 4.0cm in diameter.


1
Expert's answer
2021-12-09T09:23:35-0500

"F/A=S\cdot \Delta x/L\to \Delta x =\frac{F\cdot L}{S\cdot A}=\frac{600\cdot0.007}{1\cdot10^9\cdot(3.14\cdot0.04^2/4)}="


"=3.34\cdot10^{-6}\ (m)" . Answer